# A Longitudinal, Multinational, and Multilingual Corpus of News Coverage of the Russo-Ukrainian War

### v1

@@ 5 Use Case Experiments
    direct statements of observable or confirmed events (factual);
@@
    claims attributed to specific actors (attributed);
@@
    explicit denials or rebuttals (disputed);
  • (4)
    indirect accounts from third-party media (reported);
  • (5)
    official statements of intent or directive (command/policy); and
  • (6)
    interpretive or evaluative judgments by experts or institutions (analytical/assessment).
This approach reveals distinct narrative patterns across geopolitical boundaries: Russian sources might characterize an airstrike as precision targeting of military assets (factual/attributed), while American, British, or Ukrainian outlets describe the same event as civilian infrastructure targeting, often citing eyewitnesses (disputed/reported). Analyzing claim pairs shows entailment occurring primarily within aligned media ecosystems: Ukrainian and Western sources reinforce each other through overlapping factual, reported, or command/policy claims, while Russian media maintains internal consistency through attributed and analytical claims citing official statements. Contradictions predominantly emerge across geopolitical divides as competing interpretations rather than direct disputes. Outlets rarely offer explicit rebuttals, instead employing different topical framing, omission, or counter-attribution strategies.
Table 7 presents contradictory claim pairs from Russian and Ukrainian media. While referring to the same events, the claims construct incompatible interpretations of responsibility and action through attribution and selective emphasis.
6 Conclusion
The dnipro corpus represents a significant advancement in resources for computational analysis of media narratives during international conflicts. By integrating diverse geopolitical perspectives, consistent translations, and structured metadata across 246K articles, this dataset enables systematic study of narrative divergence that previous resources could not support. Our exploratory applications demonstrate dnipro’s utility across multiple analytical dimensions, from stance and sentiment detection to topical framing and contradiction analysis. The findings reveal how global media ecosystems construct parallel realities through attribution choices, selective emphasis, and strategic framing, rarely engaging with competing narratives.
Beyond its utility in computational journalism, dnipro provides a foundation for developing more sophisticated methods to detect cross-cultural narrative divergence, identify emergent propaganda strategies, and understand information warfare dynamics. As the global information landscape grows increasingly complex, such resources become essential for technical advancement as well as a broader understanding of narratives that shape our perception of international conflicts. limitations and ethics don’t count towards the page limit
7 Limitations
The Russo-Ukrainian War has not concluded. The dataset, which currently spans from Feb 2022 to August 2024, may omit important narratives and developments resulting from politically significant events thereafter (e.g., foreign policy changes of a powerful geopolitical entity). For the chosen news sources, a wider variety of publications representing NATO nations would be beneficial. Currently, only two of the thirty-two NATO nations are represented (the United Kingdom and the United States of America). There are other media sources like DW in Germany, which are highly referenced globally, so they are an influential factor. Additionally, representation of more countries with ambiguous involvement in the war would add value to a truly global version of the dataset. Potentially interesting countries would be Hungary and Romania, both of which border Ukraine; Belarus, also bordering Ukraine, but known to be an ally of the Russian state; Turkiye, a powerful regional actor, especially in the Black Sea; and Georgia, a NATO contender that continues to maintain reasonably strong relations with Russia.
Another facet limiting dnipro is the imperfect ground-truth annotations for stance, and other discursive phenomena. While this is not a limitation for qualitative analyses or mixed-method approaches, it does hinder nuanced supervised learning directly from the corpus. Our exploratory analyses, while illustrative, are not exhaustive and serve primarily to demonstrate the dataset’s potential rather than to make definitive empirical claims.
8 Ethical Considerations
dnipro contains potentially sensitive content related to an ongoing conflict with intense geopolitical implications. We acknowledge several potential ethical dimensions:
  • Data collection: All content was gathered from publicly available news sources without breaching access restrictions or terms of service. We do not include any copyright-protected content or any private communications.
  • Bias, representation, and inclusion: Our corpus inherently reflects the biases present in global media coverage. The inclusion of opposing narratives—including state-controlled media—is a deliberate choice, to empower comparative analysis. We strongly caution users against interpreting the mere inclusion of a source as endorsement of its content.
  • Potential misuse: While designed for media analysis, computational linguistics, and conflict studies, this resource could also be used for propaganda analysis. And thus, it could also be misused for propaganda analysis. We strongly urge the community to utilize this corpus for constructive research into understanding narrative differences to bridge societal gaps, rather than weaponize any findings.
  • Lost in translation: Machine translation introduces potential distortions, particularly with politically charged terminology and diplomatic language use. We document the details of our translation process. Further, we retain the original articles for enhanced scrutiny and possible resolution by experts in Russian, Ukrainian, and Mandarin Chinese.
  • Contextualization: We provide robust metadata (publication date, source country, outlet ownership) to ensure articles can be interpreted within their proper contexts rather than in isolation.
By openly addressing these considerations, we aim to enable valuable research on contentious discourse while minimizing potential harms. It is our expectation that through the study of such discourse, researchers will be able to bring empirical enhancements to conflict resolution mechanisms. Applications of such datasets are continuously evolving, as is the dataset itself. As such, we earnestly welcome community feedback on additional safeguards.
Appendix A Appendix
A.1 Prompt Design Details
For our quantitative analysis, we used the mistral-small-latest model via the Mistral API to jointly extract stance targets and classify stances, and employed the siebert/sentiment-roberta-large-english model from Hugging Face to estimate sentiment polarity toward named entities. We used GPT-3.5 (via the OpenAI API) to assist with annotation for three tasks: article-level tag extraction, claim identification, and claim type classification. Prompt configurations varied by task. We employed both zero-shot and few-shot strategies, along with system-level instructions and output formatting constraints, following best practices in prompt engineering for language models Saravia (2022).
A.1.1 Annotation for Stance
For stance detection, we used the mistral-small-latest model via the Mistral API. The task was to classify whether a piece of news content expressed a stance FAVOR, AGAINST, or NONE toward a generated target. We followed a combined target generation at first followed by stance classification.  Akash et al. (2025) We ran these experiments on 501 articles that mentioned the Bucha massacre dating from March to April 2022.
The following zero-shot prompt was used to guide the model:
Prompt:
Listing 1: Zero-shot prompt for joint target generation and stance classification.
Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a specified target. Analyze the following text, generate the target for this text, and determine its stance toward the generated target. A target should be the topic discussed in the text. The target can be a single word or a phrase, but its maximum length must be 4 words.
If the stance is in favor of the generated target, write FAVOR; if it is against the target, write AGAINST; if it is ambiguous, write NONE. The answer must be exactly one of these three words: FAVOR, AGAINST, or NONE.
Output format:
Target: <target>, Stance: <stance>
A.1.2 Annotation for Sentiment
We used the siebert/sentiment-roberta-large-english model from Hugging Face to perform sentence-level sentiment classification on all news articles. For each entity of interest (e.g., Russia, Ukraine, Bucha), we first extracted all sentences from the article that mentioned the entity, using spaCy’s English language model for sentence segmentation.
Prompt:
What is the sentiment(positive, negative, or neutral) towards {entity} in this context: ’{text}’?
A.1.3 Annotation for Article-level Tags
To annotate articles with high-level thematic tags, we used a zero-shot prompt with GPT-3.5 that instructed the model to extract a set of concise, specific tags summarizing each article’s content. Tags were intended to capture relevant people, locations, military units, political themes, and ideological framings. To ensure consistency and support post-processing, the prompt enforced a structured JSON output format.
Prompt:
You are a geopolitical analyst trained to extract concise topical tags from news articles related to international conflict, especially the Russia–Ukraine war. Your task is to read the content of a single article and extract 5 to 12 relevant tags that capture the article’s key topics, actors, places, events, or narrative themes. Each tag should be 1–4 words long, lowercase unless a proper noun, and specific rather than general. Avoid vague tags like “fake”, “news”, or “war” unless central to the article. Tags may include named individuals, military groups, operations, locations, events, or ideological narratives. Output the result in valid JSON using the following structure:
  "article_123": [
    "bucha incident",
    "sergei lavrov",
    "false flag allegations",
    "media manipulation",
    "un security council"
A.1.4 Annotation for Claim Extraction with Summary and Claim Types
We used a structured few-shot prompt with GPT-3.5 to extract and classify claims from news articles. Each prompt instructed the model to identify distinct claims within an article, summarize them neutrally, and assign a claim_type based on a predefined taxonomy. Claims could span one or more sentences and were returned in a structured JSON format to support post-processing.
Prompt:
You are a geopolitical analyst trained to extract claims from news articles about the Russia–Ukraine war. Your task is to read the content of a single article and extract each distinct claim it makes. For each claim, return the original sentence(s), a short neutral summary, and a claim_type. Claims may span multiple sentences. Use only the article content field for analysis. The claim_type must be one of the following: Factual, Attributed, Disputed, Reported, Command/Policy, or Analytical/Assessment. Output the result in valid JSON using the following structure:
  "article_123": {
    "claim_1": {
      "sentence":
        "Ukrainian Side Reports Multiple
        Casualties in Kyiv Airstrike.",
      "summary":
        "Ukraine reports casualties in
        Kyiv airstrike.",
      "claim_type": "Attributed"
    "claim_2": {
      "sentence":
        "Zelensky Orders: ’Ensure NATOs
        Long-Term Presence in the Black
        Sea’.",
      "summary":
        "Zelensky orders long-term NATO
        presence in the Black Sea.",
      "claim_type": "Command/Policy"
A.2 Use Case Experiment Details
Figures 4(a) and 4(b) illustrate 5-day interval favor stance percentages toward Russia and Ukraine across news source countries during March–April 2022.
(a) favor stance towards Russia (by source country).
(b) favor stance towards Ukraine (by source country).
Figure 4: 5-day interval favor stance percentages toward Russia and Ukraine (during the Bucha massacre, March–April 2022). News outlets are grouped by their national affiliations.

### v2

@@ Appendix C. Annotation Guidelines
    favor: positive stance (support, praise, endorsement);
@@
    against: negative stance (criticism, blame, opposition); or
@@
    none: No clear stance, or purely descriptive text, or ambiguous stance.
The annotators are provided these instructions:
  • Use the full passage to resolve ambiguous pronouns or implicit references.
  • If the target is none, the stance is also none.
  • Focus on the stance expressed in the sentence, not your personal opinion.
  • Implicit stances (e.g., through selective vocabulary) should be labeled.
Table 8 and Table 9 provide annotated examples from the gold-standard human-annotated validation set, illustrating the topical framing and stance annotation tasks, respectively. For stance annotation, each sentence is paired with a stance label assigned toward a target entity. For topical framing, each passage may be associated with one or more topical frames, reflecting the co-occurrence of themes within the same article. The highlighted spans indicate the textual evidence that motivated the assigned labels. These highlights are intended to make the annotation rationale explicit and to ease reader interpretation; they were not provided to human annotators, and as such, the annotations are not derived from any span-level supervision.
Passage Frames
Title: Putin gives honorary title to Russian brigade accused of war crimes in Bucha
Excerpt: A Russian brigade accused of committing war crimes in Bucha was awarded an honorary title by President Vladimir Putin. Putin congratulated the unit for its "great heroism and courage" and awarded it the title of "Guards" for "protecting Russia’s sovereignty." International leaders condemned the alleged atrocities, while investigators cited "reasonable grounds" to believe war crimes were committed.
$\bullet$ morality: 0.9
$\bullet$
politics: 0.6
$\bullet$
legal: 0.6
$\bullet$
security: 0.6
Table 8: Topical framing annotation examples showing multi-label assignments with confidence scores.
Sentence (with target) Stance
Russia has failed to capture Kyiv or Kharkiv, Ukraine’s two biggest cities, and suffered heavy losses against a far stronger Ukraine than even its allies expected. favor
Aurel Braun, a professor at the University of Toronto who has written presciently on how western governments and companies misjudged the irrationalities of the Putin regime, points out that international relations is full of theories about how the interplay of powerful corporate interests and government produces policy. against
When we see Russia’s unprovoked and unjustified invasion of Ukraine and massive disinformation campaigns and information manipulation, it is essential to separate lies from facts. against
US officials have calculated that Putin, facing crippling sanctions across other sectors of his economy, can’t afford to weaponize his energy supplies by cutting off sales to Europe. none
Vladimir Putin awarded Heroes of the Russian Federation in the Yekaterininsky Hall. none
Table 9: Stance annotations from the gold-standard human validation set, shown with highlighted target entities.